A student at a university has been doing a project to investigate entertainment habits of students. They have surveyed 100 random students who were each asked three questions: Have you watched a movie in the last week? Have you listened to music in the last week? Have you read a book in the last week? However, the student keeps a very messy room and has lost some of the results. They have been able to find the following results: Of all the students surveyed, 47 had watched a movie, 37 had listened to music and 43 had read a book in the last week. Also: 8 students answered yes to all three questions 11 students answered yA student at a university has been doing a project to investigate entertainment habits of students. They have surveyed 100 raes to questions 1 and 2 only 11 students answered yes to questions 1 and 3 only 17 students answered yes only to question 1 21 students answered no to all three questions Find the missing information and answer the following questions regarding the group surveyed. Give your answers as whole numbers. Calculate the number of students that:
Answer #1

answered we answered let A be the event that Students had watched a movie make B = be the event that students had listened CaNori nca nenes) en (Ansone) = n(B9ne). 3 21717 in (eenee :: 38 = n(BUC) 7 atas total ho. of student =1007 2 nl Bue) = 100-38.Again | MC AO BOC) + m(Ao boc) - DCAO 8) = nCanB Dee) = 11-8-3 By @@ we get ncaņene), in Canbene) In Cambo) E 3+3+ 10 =16 oc